OpenStudy (peterg2001):

Which of the following expressions has 3 terms? 8x + y - z xyz y^3 + 1

OpenStudy (anonymous):

a and b or 8x+y-z and xyz

OpenStudy (peterg2001):

@Monkey4224

OpenStudy (peterg2001):

one only

OpenStudy (calculusxy):

Terms are separated by an operation. If we look at the first choice we have : 8x + y - z which also means \( 8 \times x + y - z \) so we have here four different terms If we look at the second choice, we have xyz which also means: \(x \times y \times z\) \( \leftarrow\) now that has three terms

OpenStudy (peterg2001):

u can pick only one in her

OpenStudy (anonymous):

b seems correct

OpenStudy (anonymous):

@Monkey4224 yes she/or he is correct B is the right answer

OpenStudy (peterg2001):

i meant here